Cereal boxes are filled at a factory. Through extensive sampling, it is found that their equipment fills the boxes to a mean of 502 grams, with a standard deviation of 3.1 grams.

10,000 boxes of cereal are produced each day. 

In a daily run, how many cereal boxes are expected to weigh between 500 and 504 grams?



1
Expert's answer
2021-12-21T18:36:49-0500

"\mu=502,\sigma=3.1
\\ X\sim N(\mu,\sigma)
\\ P(500<X<504)=P(X<504)-P(X<500)
\\=P(z<\dfrac{504-502}{3.1})-P(z<\dfrac{500-502}{3.1})"

"=P(z<\dfrac{2}{3.1})-P(z<\dfrac{-2}{3.1})
\\=P(z<0.65)-P(z<-0.65)
\\=2P(z<0.65)-1
\\=2(0.74215)-1
\\=0.4843"

So, required no. of cereal boxes "=10000\times0.4843=4843"